OpenStudy (anonymous):

If f(x) = 3x − 2 and g(x) = 2x − 1, find (f − g)(2). 7 1 −7 −1

OpenStudy (anonymous):

f(2)=3(2)-2; 4 g(2)=2(2)-1; 3 therefore (f-g)(2) is (4-3) which is 1